# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

### v9

@@ 6. Unicity Problems of Algebroid Functions
Assume that is a non-compact complete Kähler manifold.
@@ 6.1. Propagation of Algebraic Dependence
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Put
@@
Let be an effective divisor on and fix any positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
It is clear that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and be positive integers (allowed to be ).
We introduce the notation as follows.
When is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), denote by
@@
When has non-positive sectional curvature satisfying (3), denote by
the set of all such that
and (5) is satisfied.
Let be the hyperplane line bundle over It gives a holomorphic line bundle over
where is the natural projection on the -th factor with Let for For , define
@@
where is a fixed positive -form.
@@
Apply Dynkin formula, we deduce that
@@
Let be the set of all indecomposable hypersurfaces in such that for some Put
@@
By we obtain
@@
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the hyperplane line bundle over If on then there is a holomorphic section of over such that and Taking then we obtain and On the other hand, implies that
@@ 6.2. Five-Value Type Theorem
in the sense of currents, where and are understood as divisors. So, we are led to
It yields from Lemma 6.3 that
Similarly,
Combining (6.2) with (6.2), it yields from Theorem 5.4 that

### v10

@@ 6. Unicity Problems of Algebroid Functions
Assume that is a complete non-compact Kähler manifold.
@@ 6.1. Propagation of Algebraic Dependence
Fix any integer A proper algebraic subset of is said to be decomposible, if there exist positive integers with for some integer and algebraic subsets with such that If is not decomposable, then we say that is indecomposable. Given algebroid functions Set
@@
We let be an effective divisor on and fix a positive integer (allowed to be ). Write for distinct irreducible hypersurfaces of and positive integers Define the support of with order at most by
Note that Set
Let be hypersurfaces of satisfying that if or if for Let be distinct values in and let be positive integers (allowed to be ).
We introduce the notation as follows:
If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then denote by
@@
Let be the point line bundle over It gives a holomorphic line bundle over :
where is the natural projection on the -th factor with Let with For , define
@@
where is a positive -form.
@@
Applying Dynkin formula, we deduce that
@@
We denote by the set of all indecomposable hypersurfaces of such that for some Set
@@
Since we obtain
in the sense of currents, where and are understood as divisors. Thus, we are led to
@@
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the point line bundle over If on then there is a holomorphic section of over such that and Taking then one can have and On the other hand, implies that
@@ 6.2. Five-Value Type Theorem
In further, it follows from Lemma 6.3 that
Similarly, we have
Combining the above and using Theorem 5.4, we are led to